Group Selection Page HELP Controls Group ITDM Vacuum WP78 FXE SPB HED SQS SXP Sample Environment PSPO XO TS Migrated2Zulip Archived2GitlabPages Archived2PDF MigratedByGroup
SCS Photon Commissioning Team XRO DOC WP74 MID XO_mig XO_Operation SQS Migrated2Zulip DET
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2024 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn  Not logged in ELOG logo
Message ID: 160     Entry time: 09 Sep 2024, 06:52
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 12.8 keV, 1 pulse, ~0.7 mJ

Achievements/Observations

  • Except for some downtime due to beam down, majority of the time was for data collection
  • 6 windows of a chip collected, with variations in shots spacing (30 um and 60 um) and rotation

Issues

  • Something is still causing beam drift constantly, even though we can correct it with the feedback.
ELOG V3.1.4-7c3fd00